OpenStudy (anonymous):

Evaluate 2x2 − 5y when x = -3 and y = 4

OpenStudy (anonymous):

(8x+3y)/5y =(56+12)/20 =68/20

OpenStudy (anonymous):

not one of the answers

sammixboo (sammixboo):

Hi @BigBoy6191 Welcome to \(OpenStudy!\) \(68/20\) can equal many things such as: ~ \(3\frac{8}{20}\) ~ \(3\frac{2}{5}\) ~ \(\frac{17}{5}\) ~ \(\frac{34}{10}\) ^^If any of those numbers above is in your answer choices then it is correct!^^ If none of those numbers are not in your choices then just tell me! If you hover your mouse over a person's username and click it, and a box will appear, you can view their SmartScore, send them messages and become there fan!